# FunAudio-ASR Technical Report

### v2

@@ Abstract
In recent years, automatic speech recognition (ASR) has witnessed transformative advancements driven by three complementary paradigms: data scaling, model size scaling, and deep integration with large language models (LLMs). However, LLMs are prone to hallucination, which can significantly degrade user experience in real-world ASR applications. In this paper, we present FunAudio-ASR, a large-scale, LLM-based ASR system that synergistically combines massive data, large model capacity, LLM integration, and reinforcement learning to achieve state-of-the-art performance across diverse and complex speech recognition scenarios. Moreover, FunAudio-ASR is specifically optimized for practical deployment, with enhancements in streaming capability, noise robustness, code-switching, hotword customization, and satisfying other real-world application requirements. Experimental results show that while most LLM-based ASR systems achieve strong performance on open-source benchmarks, they often underperform on real industry evaluation sets. Thanks to production-oriented optimizations, FunAudio-ASR achieves SOTA performance on real application datasets, demonstrating its effectiveness and robustness in practical settings.
Figure 1: Performance comparison (in Accuracy) between our FunAudio-ASR (7.7B) and FunAudio-ASR-nano (0.8B) and top-tier ASR and speech-text multimodal models, including Paraformer-v2, Kimi-Audio (8B), FireRedASR (1.1B), and Seed-ASR ($\sim$12B).
@@ 1 Introduction
Building upon these significant advancements, we propose FunAudio-ASR, a large-scale LLM-based ASR system trained on large-scale data. FunAudio-ASR exhibits the following key characteristics:
  • Scaling and Innovative LLM Integration. FunAudio-ASR is designed to harness the synergistic benefits of data scaling, model size scaling, and LLM integration.
  • State-of-the-art speech recognition accuracy. Through synergistic advancements in data scaling, model size scaling, and innovative architectural integration with LLMs, FunAudio-ASR achieves unprecedented recognition accuracy across diverse linguistic and acoustic domains, establishing a new state of the art for ASR systems. Our comprehensive evaluations demonstrate that FunAudio-ASR substantially outperforms both our previous small-scale models and leading ASR systems in industry, in terms of critical metrics across multiple challenging benchmarks.
  • Optimization for practical production usage. Beyond achieving state-of-the-art performance on standardized benchmarks, FunAudio-ASR is meticulously engineered to meet the complex demands of real-world deployment scenarios, with a particular focus on practical usability, reliability, and user experience. We implement a comprehensive suite of optimizations in multiple dimensions, each addressing specific challenges encountered in commercial applications. (1) First, we implement a highly efficient streaming ASR architecture for FunAudio-ASR that supports real-time processing with minimal latency, enabling seamless integration into live applications such as video conferencing, live captioning, and voice-controlled devices. (2) Second, we enhance noise robustness substantially through a multi-stage approach. (3) Third, we implement advanced code-switching capabilities that seamlessly handle transitions between Chinese and English within the same utterance, which is critical for multilingual users in global business environments. (4) Fourth, we integrate customizable hotword recognition that allows users to define domain-specific terms or phrases for enhanced recognition accuracy. This feature is particularly valuable in specialized domains such as healthcare, enterprise, and automotive technology.
    To thoroughly evaluate these production-oriented optimizations, we develop a comprehensive evaluation protocol that includes both standardized benchmarks and real-world usage scenarios. This protocol encompasses distinct test sets, each simulating different application contexts. Our evaluation results reveal that FunAudio-ASR not only excels in recognition accuracy but also provides superior practical performance. These results demonstrate that FunAudio-ASR successfully bridges the gap between academic research and commercial production readiness, offering a comprehensive solution to address real-world speech recognition challenges.
@@ 2 Model Architecture
Figure 2: Overview of the FunAudio-ASR model architecture.
FunAudio-ASR comprises four key components: (1) an audio encoder that extracts representations from the input speech, implemented as multiple layers of transformer encoder, (2) an audio adaptor that connects the audio encoder output with the LLM, implemented as two layers of transformer encoder, (3) a CTC decoder that is built upon the audio encoder to obtain the initial recognition hypothesis, which will be used for hotword customization, as described in Section 5.5, and (4) an LLM-based decoder that produces output based on the audio condition and the CTC prediction.
To address varying constraints on computational resources and requirements on inference efficiency, we propose two models with different model sizes: FunAudio-ASR and FunAudio-ASR-nano. FunAudio-ASR comprises an audio encoder with 0.7B parameters and an LLM-based decoder with 7B parameters, aiming for high recognition accuracy; FunAudio-ASR-nano comprises an audio encoder with 0.2B parameters and an LLM-based decoder with 0.6B parameters, seeking to strike a balance between accuracy and efficiency to meet the demands of low-computational-resource scenarios.
@@ 4.3 Contextual Supervised Fine-tuning
As a content prior, context can effectively help the model identify and disambiguate key text content from easily confusable pronunciation in ASR tasks, and improve the accuracy of long-term continuous recognition in complex scenarios. Consequently, after the SFT training (Section 4.2), we further train FunAudio-ASR on the contextual and long-duration data to enhance its contextual modeling capability.
@@ 4.4.1 The RL Framework for Large Audio-Language Models
We design FunRL, an efficient reinforcement learning (RL) framework tailored for large audio-language models (LALMs). Different from text LLMs, FunAudio-ASR, as an LALM, incorporates an audio encoder to convert speech into embeddings—a component that is not natively supported by existing RL frameworks such as Verl (Sheng et al., 2024) or Trl (von Werra et al., 2020). As illustrated in Figure 4(a), FunRL orchestrates the audio encoder, rollout, and policy modules using Ray, enabling them to alternately utilize GPU resources. In the audio encoder inference stage, all input audio clips are batched and processed through a Torch-based encoder. The encoder extracts audio embeddings in parallel and transfers the resulting embeddings from GPU to CPU. Subsequently, the SGLang-based LLM rollout takes control of the GPU to generate multiple hypothesis sequences based on the audio embeddings and the instruction text tokens. Each hypothesis is assigned a reward according to the predefined rules, which will be detailed later. Finally, the FSDP-based LLM policy model uses the audio embeddings and the generated hypotheses to compute output probabilities and performs policy optimization via RL. After each update, the optimized policy is synchronized back to the rollout module, ensuring that the RL process remains on-policy.
@@ 4.4.2 GRPO-based RL for ASR
Based on the FunRL framework, we enhance the GRPO-based RL algorithm for FunAudio-ASR. Among various RL algorithms, GRPO is a lightweight and effective policy-based method that has achieved significant success in LLM post-training. Unlike RL algorithms such as PPO, GRPO generates a group of responses $\{o_{i}\}_{i=1}^{G}$ and computes a rule-based value function to assign a reward $\{R_{i}\}_{i=1}^{G}$ to each response. These group-level rewards are then normalized to compute the advantage $\hat{A}_{i,t}$, which guides the policy update.
@@ 4.4.3 Constructing the RL Training Data
  • Hardcase Samples. We collect a large amount of unlabeled speech and transcribe each utterance using the FunAudio-ASR model after contextual SFT (referred to as the base model), along with three other distinct ASR systems, including Whisper, FireRed-ASR, and SenseVoice. If the outputs from the three external systems are consistent with each other (WER < 5%) but differ significantly from FunAudio-ASR’s output (WER > 10%), the sample is identified as a hard case and is included in the RL training set.
@@
The final RL training data consist of 100K samples, with 20K utterances in each of the five subsets described above. Thanks to the efficiency of the FunRL framework, the entire RL training of FunAudio-ASR can be completed within one day with 8 A100 GPUs.
@@ 5.1 Streaming Ability
To enhance the streaming capability of large audio language model FunAudio-ASR, we construct streaming-style training data that explicitly emulate the streaming decoding process, thereby reducing the mismatch between training and inference. Specifically, we sample a subset of the offline training corpus and transform it into incremental, chunked inputs that expose only past context. Fine-tuning by combining this simulated streaming with previous offline training data improves the model’s performance under streaming decoding.
@@ 5.2 Noise Robust Training
Given the diverse range of real-world deployment scenarios, it is essential for FunAudio-ASR to maintain reliable performance under challenging acoustic conditions, such as those found in restaurants, train stations, and shopping malls, without significant performance degradation. However, creating a dataset that fully captures the complexity and variability of real-world noise environment is impractical. To address this challenge, we employ a large-scale noisy data augmentation strategy. We begin by selecting approximately 110K hours of low-noise speech and 10K hours of noise samples from our in-house corpus. These are combined to generate around 110K hours of offline simulated noisy speech, with an average signal-to-noise ratio (SNR) of 10 dB and a standard deviation of 5 dB. To further increase data diversity, 30% of the training speech is randomly chosen for online data augmentation, where environmental noise is mixed during training. This comprehensive approach for noise robustness yields approximately 13% average relative performance improvement on our complex-noise evaluation set.
@@ 5.3 Multilingual ASR
The availability of training data varies widely across different languages. Resource-rich languages such as Chinese and English have abundant data, whereas languages such as Vietnamese and Thai have comparatively limited resources. The primary FunAudio-ASR model is a Chinese-English model. To improve multilingual ASR performance, we train an additional multilingual FunAudio-ASR model to support these languages. The current multilingual FunAudio-ASR model (denoted by FunAudio-ASR-ML) supports Chinese, English, Vietnamese, Thai, and Indonesian, with a broader language coverage planned for future releases. During training, we downsample the Chinese and English data used for the Chinese-English FunAudio-ASR model and upsample the Vietnamese, Thai, and Indonesian data to balance the distribution. In total, this multilingual dataset comprises approximately 500K hours of audio. The training methodology is the same as that used for the Chinese-English FunAudio-ASR model (Section 4).
@@ 5.5 Hotword Customization
Within FunAudio-ASR, we implement a RAG-based mechanism for hotword customization. Specifically, we construct a hotword vocabulary in which each prespecified hotword is converted into a phoneme sequence (for Chinese) or a word-piece sequence (for other languages) using a predefined lexicon. During inference, we retrieve hotword candidates from the vocabulary based on the phoneme-level or the word-piece-level edit distance between the CTC hypotheses and entries in the hotword vocabulary. The retrieved hotword candidates, together with the audio input and the CTC prediction, are used as the input to the LLM, as depicted in Figure 2, to produce the hotword-customized output.
@@ 5.6 Hallucination Mitigation
Hallucination in ASR, where an ASR system generates text that is not present in the input audio, is particularly problematic during silence, abrupt speaker interruptions, or in noisy environments where the model may produce spurious transcriptions even without speech. To mitigate hallucination, FunAudio-ASR adopts the following strategy. During data augmentation, we introduce zero-padding into audio signals before adding noise, thereby creating pure-noise segments. This strategy forces the model to learn recognizing noise-only inputs and aligning its outputs accordingly, hence reducing the likelihood of hallucinated text. We find that this approach helps enhance the robustness, accuracy, and stability of FunAudio-ASR across diverse acoustic conditions.
@@ 6.1 Evaluation Setting
We evaluate FunAudio-ASR and FunAudio-ASR-ML on both open-source ASR benchmark datasets and real-world industry evaluation sets. For the open-source evaluation, we use corresponding test sets of AIShell-1 (Bu et al., 2017), AIShell-2(Du et al., 2018), Librispeech(Panayotov et al., 2015), Fleurs(Conneau et al., 2022), WeNetSpeech(Zhang et al., 2022), Gigaspeech2 (Yang et al., 2025c) data sets. These open-source datasets have been publicly available for a long time, increasing the risk of data leakage into model training sets. To ensure a more reliable and leakage-free evaluation, we collect newly uploaded videos from YouTube and Bilibili posted after June 30th, which are then manually transcribed to form an independent test set. For noise robustness evaluation, we use real-world audio recordings captured in various environments, including canteen, dinner, meeting, office, outdoor, park, shop, street, subway, supermarket, and walk-street. These are further categorized by acoustic conditions and topic to better assess performance under diverse and challenging scenarios.
@@ 6.2.1 Overall Results
We first evaluate recently published ASR systems on open-source benchmarks, with results shown in Table 1. On these datasets, all models achieve very low WER, and some open-source models even outperform commercial APIs on Librispeech and AIShell. However, as shown in Table 2, on real industry evaluation sets, Seed-ASR-API demonstrates a clear advantage over other open-source models, particularly in noisy conditions. This indicates that performance on open-source test data may not reliably reflect real-world ASR capabilities, highlighting the importance of regularly updating evaluation sets to prevent data leakage. Compared to both open-source models and commercial APIs, our FunAudio-ASR achieves SOTA performance on both open-source benchmarks (Table 1) and industry datasets (Table 2). Since all training data is collected before June 30th 2025, we ensure no data leakage during our evaluations, making the results trustworthy and reproducible. Notably, FunAudio-ASR-nano also outperforms the open-source models and performs closely to Seed-ASR, with only 0.8B parameters.
Whisper- Seed- Seed- Kimi- Step- FireRed- FunAudio- FunAudio- Test set large-v3 ASR ASR* Audio Audio2 ASR ASR-nano ASR AIShell1 4.72 0.68 1.63 0.71 0.63 0.54 1.8 1.22 AIShell2 4.68 2.27 2.76 2.86 2.10 2.58 2.95 2.3 Fleurs-zh 5.18 3.43 3.23 3.11 2.68 4.81 3.47 2.64 Fleurs-en 6.23 9.39 9.39 6.99 3.03 10.79 8.42 5.84 Librispeech-clean 1.86 1.58 2.8 1.32 1.17 1.84 1.94 1.57 Librispeech-other 3.43 2.84 5.69 2.63 2.42 4.52 4.69 3.24 WenetSpeech Meeting 18.39 5.69 7.07 6.24 4.75 4.95 6.82 6.49 WenetSpeech Net 11.89 4.66 4.84 6.45 4.67 4.94 6.04 5.46
@@
Seed- Whisper- FireRed- Kimi- Paraformer FunAudio- FunAudio- Test set ASR large-v3 ASR Audio v2 ASR-nano ASR In-house 7.20 16.58 10.10 9.02 8.11 7.89 6.66 Fairfield 4.59 22.21 7.49 10.95 9.55 5.75 4.66 Home Scenario 8.08 18.17 9.67 23.79 6.87 6.94 5.17 Complex Background 12.90 32.57 15.56 15.56 15.19 16.17 11.29 English General 15.65 18.56 21.62 18.12 19.48 16.34 14.22 Opensouce 3.83 7.05 5.31 3.20 6.23 4.52 3.60 Average 8.71 19.19 11.63 13.54 10.91 9.60 7.60

### v3

# Fun-ASR Technical Report

@@ Abstract
In recent years, automatic speech recognition (ASR) has witnessed transformative advancements driven by three complementary paradigms: data scaling, model size scaling, and deep integration with large language models (LLMs). However, LLMs are prone to hallucination, which can significantly degrade user experience in real-world ASR applications. In this paper, we present Fun-ASR, a large-scale, LLM-based ASR system that synergistically combines massive data, large model capacity, LLM integration, and reinforcement learning to achieve state-of-the-art performance across diverse and complex speech recognition scenarios. Moreover, Fun-ASR is specifically optimized for practical deployment, with enhancements in streaming capability, noise robustness, code-switching, hotword customization, and satisfying other real-world application requirements. Experimental results show that while most LLM-based ASR systems achieve strong performance on open-source benchmarks, they often underperform on real industry evaluation sets. Thanks to production-oriented optimizations, Fun-ASR achieves SOTA performance on real application datasets, demonstrating its effectiveness and robustness in practical settings.
Figure 1: Performance comparison (in Accuracy) between our Fun-ASR (7.7B) and Fun-ASR-nano (0.8B) and top-tier ASR and speech-text multimodal models, including Paraformer-v2, Kimi-Audio (8B), FireRedASR (1.1B), and Seed-ASR ($\sim$12B).
@@ 1 Introduction
Building upon these significant advancements, we propose Fun-ASR, a large-scale LLM-based ASR system trained on large-scale data. Fun-ASR exhibits the following key characteristics:
  • Scaling and Innovative LLM Integration. Fun-ASR is designed to harness the synergistic benefits of data scaling, model size scaling, and LLM integration.
  • State-of-the-art speech recognition accuracy. Through synergistic advancements in data scaling, model size scaling, and innovative architectural integration with LLMs, Fun-ASR achieves unprecedented recognition accuracy across diverse linguistic and acoustic domains, establishing a new state of the art for ASR systems. Our comprehensive evaluations demonstrate that Fun-ASR substantially outperforms both our previous small-scale models and leading ASR systems in industry, in terms of critical metrics across multiple challenging benchmarks.
  • Optimization for practical production usage. Beyond achieving state-of-the-art performance on standardized benchmarks, Fun-ASR is meticulously engineered to meet the complex demands of real-world deployment scenarios, with a particular focus on practical usability, reliability, and user experience. We implement a comprehensive suite of optimizations in multiple dimensions, each addressing specific challenges encountered in commercial applications. (1) First, we implement a highly efficient streaming ASR architecture for Fun-ASR that supports real-time processing with minimal latency, enabling seamless integration into live applications such as video conferencing, live captioning, and voice-controlled devices. (2) Second, we enhance noise robustness substantially through a multi-stage approach. (3) Third, we implement advanced code-switching capabilities that seamlessly handle transitions between Chinese and English within the same utterance, which is critical for multilingual users in global business environments. (4) Fourth, we integrate customizable hotword recognition that allows users to define domain-specific terms or phrases for enhanced recognition accuracy. This feature is particularly valuable in specialized domains such as healthcare, enterprise, and automotive technology.
    To thoroughly evaluate these production-oriented optimizations, we develop a comprehensive evaluation protocol that includes both standardized benchmarks and real-world usage scenarios. This protocol encompasses distinct test sets, each simulating different application contexts. Our evaluation results reveal that Fun-ASR not only excels in recognition accuracy but also provides superior practical performance. These results demonstrate that Fun-ASR successfully bridges the gap between academic research and commercial production readiness, offering a comprehensive solution to address real-world speech recognition challenges.
@@ 2 Model Architecture
Figure 2: Overview of the Fun-ASR model architecture.
Fun-ASR comprises four key components: (1) an audio encoder that extracts representations from the input speech, implemented as multiple layers of transformer encoder, (2) an audio adaptor that connects the audio encoder output with the LLM, implemented as two layers of transformer encoder, (3) a CTC decoder that is built upon the audio encoder to obtain the initial recognition hypothesis, which will be used for hotword customization, as described in Section 5.5, and (4) an LLM-based decoder that produces output based on the audio condition and the CTC prediction.
To address varying constraints on computational resources and requirements on inference efficiency, we propose two models with different model sizes: Fun-ASR and Fun-ASR-nano. Fun-ASR comprises an audio encoder with 0.7B parameters and an LLM-based decoder with 7B parameters, aiming for high recognition accuracy; Fun-ASR-nano comprises an audio encoder with 0.2B parameters and an LLM-based decoder with 0.6B parameters, seeking to strike a balance between accuracy and efficiency to meet the demands of low-computational-resource scenarios.
@@ 4.3 Contextual Supervised Fine-tuning
As a content prior, context can effectively help the model identify and disambiguate key text content from easily confusable pronunciation in ASR tasks, and improve the accuracy of long-term continuous recognition in complex scenarios. Consequently, after the SFT training (Section 4.2), we further train Fun-ASR on the contextual and long-duration data to enhance its contextual modeling capability.
@@ 4.4.1 The RL Framework for Large Audio-Language Models
We design FunRL, an efficient reinforcement learning (RL) framework tailored for large audio-language models (LALMs). Different from text LLMs, Fun-ASR, as an LALM, incorporates an audio encoder to convert speech into embeddings—a component that is not natively supported by existing RL frameworks such as Verl (Sheng et al., 2024) or Trl (von Werra et al., 2020). As illustrated in Figure 4(a), FunRL orchestrates the audio encoder, rollout, and policy modules using Ray, enabling them to alternately utilize GPU resources. In the audio encoder inference stage, all input audio clips are batched and processed through a Torch-based encoder. The encoder extracts audio embeddings in parallel and transfers the resulting embeddings from GPU to CPU. Subsequently, the SGLang-based LLM rollout takes control of the GPU to generate multiple hypothesis sequences based on the audio embeddings and the instruction text tokens. Each hypothesis is assigned a reward according to the predefined rules, which will be detailed later. Finally, the FSDP-based LLM policy model uses the audio embeddings and the generated hypotheses to compute output probabilities and performs policy optimization via RL. After each update, the optimized policy is synchronized back to the rollout module, ensuring that the RL process remains on-policy.
@@ 4.4.2 GRPO-based RL for ASR
Based on the FunRL framework, we enhance the GRPO-based RL algorithm for Fun-ASR. Among various RL algorithms, GRPO is a lightweight and effective policy-based method that has achieved significant success in LLM post-training. Unlike RL algorithms such as PPO, GRPO generates a group of responses $\{o_{i}\}_{i=1}^{G}$ and computes a rule-based value function to assign a reward $\{R_{i}\}_{i=1}^{G}$ to each response. These group-level rewards are then normalized to compute the advantage $\hat{A}_{i,t}$, which guides the policy update.
@@ 4.4.3 Constructing the RL Training Data
  • Hardcase Samples. We collect a large amount of unlabeled speech and transcribe each utterance using the Fun-ASR model after contextual SFT (referred to as the base model), along with three other distinct ASR systems, including Whisper, FireRed-ASR, and SenseVoice. If the outputs from the three external systems are consistent with each other (WER < 5%) but differ significantly from Fun-ASR’s output (WER > 10%), the sample is identified as a hard case and is included in the RL training set.
@@
The final RL training data consist of 100K samples, with 20K utterances in each of the five subsets described above. Thanks to the efficiency of the FunRL framework, the entire RL training of Fun-ASR can be completed within one day with 8 A100 GPUs.
@@ 5.1 Streaming Ability
To enhance the streaming capability of large audio language model Fun-ASR, we construct streaming-style training data that explicitly emulate the streaming decoding process, thereby reducing the mismatch between training and inference. Specifically, we sample a subset of the offline training corpus and transform it into incremental, chunked inputs that expose only past context. Fine-tuning by combining this simulated streaming with previous offline training data improves the model’s performance under streaming decoding.
@@ 5.2 Noise Robust Training
Given the diverse range of real-world deployment scenarios, it is essential for Fun-ASR to maintain reliable performance under challenging acoustic conditions, such as those found in restaurants, train stations, and shopping malls, without significant performance degradation. However, creating a dataset that fully captures the complexity and variability of real-world noise environment is impractical. To address this challenge, we employ a large-scale noisy data augmentation strategy. We begin by selecting approximately 110K hours of low-noise speech and 10K hours of noise samples from our in-house corpus. These are combined to generate around 110K hours of offline simulated noisy speech, with an average signal-to-noise ratio (SNR) of 10 dB and a standard deviation of 5 dB. To further increase data diversity, 30% of the training speech is randomly chosen for online data augmentation, where environmental noise is mixed during training. This comprehensive approach for noise robustness yields approximately 13% average relative performance improvement on our complex-noise evaluation set.
@@ 5.3 Multilingual ASR
The availability of training data varies widely across different languages. Resource-rich languages such as Chinese and English have abundant data, whereas languages such as Vietnamese and Thai have comparatively limited resources. The primary Fun-ASR model is a Chinese-English model. To improve multilingual ASR performance, we train an additional multilingual Fun-ASR model to support these languages. The current multilingual Fun-ASR model (denoted by Fun-ASR-ML) supports Chinese, English, Vietnamese, Thai, and Indonesian, with a broader language coverage planned for future releases. During training, we downsample the Chinese and English data used for the Chinese-English Fun-ASR model and upsample the Vietnamese, Thai, and Indonesian data to balance the distribution. In total, this multilingual dataset comprises approximately 500K hours of audio. The training methodology is the same as that used for the Chinese-English Fun-ASR model (Section 4).
@@ 5.5 Hotword Customization
Within Fun-ASR, we implement a RAG-based mechanism for hotword customization. Specifically, we construct a hotword vocabulary in which each prespecified hotword is converted into a phoneme sequence (for Chinese) or a word-piece sequence (for other languages) using a predefined lexicon. During inference, we retrieve hotword candidates from the vocabulary based on the phoneme-level or the word-piece-level edit distance between the CTC hypotheses and entries in the hotword vocabulary. The retrieved hotword candidates, together with the audio input and the CTC prediction, are used as the input to the LLM, as depicted in Figure 2, to produce the hotword-customized output.
@@ 5.6 Hallucination Mitigation
Hallucination in ASR, where an ASR system generates text that is not present in the input audio, is particularly problematic during silence, abrupt speaker interruptions, or in noisy environments where the model may produce spurious transcriptions even without speech. To mitigate hallucination, Fun-ASR adopts the following strategy. During data augmentation, we introduce zero-padding into audio signals before adding noise, thereby creating pure-noise segments. This strategy forces the model to learn recognizing noise-only inputs and aligning its outputs accordingly, hence reducing the likelihood of hallucinated text. We find that this approach helps enhance the robustness, accuracy, and stability of Fun-ASR across diverse acoustic conditions.
@@ 6.1 Evaluation Setting
We evaluate Fun-ASR and Fun-ASR-ML on both open-source ASR benchmark datasets and real-world industry evaluation sets. For the open-source evaluation, we use corresponding test sets of AIShell-1 (Bu et al., 2017), AIShell-2(Du et al., 2018), Librispeech(Panayotov et al., 2015), Fleurs(Conneau et al., 2022), WeNetSpeech(Zhang et al., 2022), Gigaspeech2 (Yang et al., 2025c) data sets. These open-source datasets have been publicly available for a long time, increasing the risk of data leakage into model training sets. To ensure a more reliable and leakage-free evaluation, we collect newly uploaded videos from YouTube and Bilibili posted after June 30th, which are then manually transcribed to form an independent test set. For noise robustness evaluation, we use real-world audio recordings captured in various environments, including canteen, dinner, meeting, office, outdoor, park, shop, street, subway, supermarket, and walk-street. These are further categorized by acoustic conditions and topic to better assess performance under diverse and challenging scenarios.
@@ 6.2.1 Overall Results
We first evaluate recently published ASR systems on open-source benchmarks, with results shown in Table 1. On these datasets, all models achieve very low WER, and some open-source models even outperform commercial APIs on Librispeech and AIShell. However, as shown in Table 2, on real industry evaluation sets, Seed-ASR-API demonstrates a clear advantage over other open-source models, particularly in noisy conditions. This indicates that performance on open-source test data may not reliably reflect real-world ASR capabilities, highlighting the importance of regularly updating evaluation sets to prevent data leakage. Compared to both open-source models and commercial APIs, our Fun-ASR achieves SOTA performance on both open-source benchmarks (Table 1) and industry datasets (Table 2). Since all training data is collected before June 30th 2025, we ensure no data leakage during our evaluations, making the results trustworthy and reproducible. Notably, Fun-ASR-nano also outperforms the open-source models and performs closely to Seed-ASR, with only 0.8B parameters.
Whisper- Seed- Seed- Kimi- Step- FireRed- Fun-ASR Fun- Test set large-v3 ASR ASR* Audio Audio2 ASR -nano ASR AIShell1 4.72 0.68 1.63 0.71 0.63 0.54 1.76 1.22 AIShell2 4.68 2.27 2.76 2.86 2.10 2.58 2.80 2.3 Fleurs-zh 5.18 3.43 3.23 3.11 2.68 4.81 3.47 2.64 Fleurs-en 6.23 9.39 9.39 6.99 3.03 10.79 7.95 5.84 Librispeech-clean 1.86 1.58 2.8 1.32 1.17 1.84 1.75 1.57 Librispeech-other 3.43 2.84 5.69 2.63 2.42 4.52 4.37 3.24 WenetSpeech Meeting 18.39 5.69 7.07 6.24 4.75 4.95 8.78 6.49 WenetSpeech Net 11.89 4.66 4.84 6.45 4.67 4.94 6.28 5.46
@@
Seed- Whisper- FireRed- Kimi- Paraformer Fun-ASR Fun- Test set ASR large-v3 ASR Audio v2 -nano ASR In-house 7.20 16.58 10.10 9.02 8.11 7.26 6.66 Fairfield 4.59 22.21 7.49 10.95 9.55 5.43 4.66 Home Scenario 8.08 18.17 9.67 23.79 6.87 6.02 5.17 Complex Background 12.90 32.57 15.56 15.56 15.19 17.07 11.29 English General 15.65 18.56 21.62 18.12 19.48 15.87 14.22 Opensouce 3.83 7.05 5.31 3.20 6.23 4.65 3.60 Average 8.71 19.19 11.63 13.54 10.91 9.38 7.60